OpenStudy (anonymous):

Two electric bulbs whose resistances are in the ratio of 1 : 2 are connected in serie. The powers dissipated in them have what ratio?

OpenStudy (anonymous):

P=I\[I{2 timesR}\] so we get 1:2

OpenStudy (anonymous):

the real answer is 2:1. I need the woking

OpenStudy (anonymous):

Let bulb 1 have resistance \(R\) and bulb 2 have resistance \(2R\). The power is thus given as follows. Note that since they are in series, the current through them is going be the same.\[P_1 = I^2 R\]\[P_2 = I^2 (2R) = 2(I^2R)\]The ratio is thus 1:2. Not sure where you're getting 2:1 from; that is incorrect.

OpenStudy (anonymous):

what if they are asking the ratio of energy, would they give the ratio 2:1, if yes how?

OpenStudy (anonymous):